### v1

# MTDT: A Multi-Task Deep Learning Digital Twin

Abstract
Traffic congestion has significant impacts on both the economy and the environment. Measures of Effectiveness (MOEs) have long been the standard for evaluating the level of service and operational efficiency of traffic intersections. However, the scarcity of traditional high-resolution loop detector data (ATSPM) presents challenges in accurately measuring MOEs or capturing the intricate temporospatial characteristics inherent in urban intersection traffic. In response to this challenge, we have introduced the Multi-Task Deep Learning Digital Twin (MTDT) as a solution for multifaceted and precise intersection traffic flow simulation. MTDT enables accurate, fine-grained estimation of loop detector waveform time series for each lane of movement, alongside successful estimation of several MOEs for each lane group associated with a traffic phase concurrently and for all approaches of an arbitrary urban intersection. Compared to existing deep learning methodologies, MTDT distinguishes itself through its adaptability to local temporal and spatial features, such as signal timing plans, intersection topology, driving behaviors, and turning movement counts. While maintaining a straightforward design, our model emphasizes the advantages of multi-task learning in traffic modeling. Although additional complexities could potentially enhance feature learning, we have prioritized simplicity to underscore the benefits of multi-task learning in modeling traffic flow dynamics. By consolidating the learning process across multiple tasks, MTDT demonstrates reduced overfitting, increased efficiency, and enhanced effectiveness through the sharing of representations learned by different tasks. Furthermore, our approach facilitates sequential computation and lends itself to complete parallelization through GPU implementation. This not only streamlines the computational process but also enhances scalability and performance.
Index Terms:
Deep Learning, Multi-task Learning, Graph Neural Networks, Convolutional Neural Networks, Traffic, Intersection, ATSPM, MOEs.
I Introduction
Traffic congestion in urban areas impacts efficiency, economics, environment, safety, and overall life quality. Measures of Effectiveness (MOEs) provide a benchmark of longitudinal perspectives to quantify the level of service and evaluate the performance of transportation systems. The employment of MOEs is also useful for optimizing the signal timing plan of intersections and the overall traffic flow accordingly. Common MOEs for urban traffic control include factors such as queue length, travel time, waiting time, number of stops, vehicle throughput, average speed, fuel efficiency and emissions, safety, equity, adaptability to demand, public satisfaction, accessibility, etc. Depending on the specific municipal goals and priorities, a combination of MOEs can be used to effectively evaluate the urban traffic control system. This study includes but is not limited to two commonly used MOEs: 1- Queue Length: Distance from the stop bar to the tail of the last vehicle stopped in any traffic lane, and 2- Travel Time: The time necessary to traverse any route connects any two points of interest.
Microscopic simulators e.g., VISSIM, SUMO, and AIMSUN fail to generalize to arbitrary topology and infrastructure of traffic urban roadways without painstakingly (re)drawing base maps and (re)setting traffic simulation parameters. Some data sources e.g., GPS, Video, BlueTooth, DSRC etc. are normally sparse with privacy and ownership restrictions [19]. Advanced sensing technologies such as induction loop detectors (ATSPM) make high-resolution loop detector traffic data widely available, although it fails to provide enough information for MOEs estimation by itself. The computation of MOEs is specific to several tempo-spacial factors such as intersection characteristics(e.g., topology, lane configuration), traffic pattern (e.g., inflow traffic and turning-movement counts, signal timings, and traffic conditions (e.g., driving behaviors).
Existing analytical methods of estimating MOEs from ATSPM either use vehicle conservation input-output algorithms based on counting vehicle flow over advance and stop-bar detectors [18] or Lighthill’s shockwave algorithms based on identifying breakpoints in the detector actuation waveforms [22]. Although, both of them are sensitive to the placement of detectors and limited to the localized training datasets. Machine learning methods use either GPS trace data [5] that are too coarse in recording dissipation trends, or ATSPM for cycle-wise MOEs which is not based on real-world traffic conditions due to the random vehicle flow and road networks without explicit left-turn buffers. Recently, some deep learning methods have employed geometric deep learning [20] and graph convolution networks [7] to extract cycle-wise MOEs, although these approaches still fail to generalize the local traffic conditions and road geometries, also their reconstructed MOEs have as low resolution as approach level.
We propose a Multi-Task Deep Learning-based (MTDT) digital twin of intersections to address the aforementioned shortcomings. This model is trained on real-world data obtained from high-resolution induction loop detectors (ATSPM), capturing varying traffic patterns encompassing a wide range of configurable parameters, thus generating potentially viable counterfactual traffic scenarios. MTDT builds upon the previous work by Yousefzadeh et al. [23], extending it to create an accurate and reliable multi-purpose digital twin. An overview of our model is depicted in Figure 1, where orange-colored objects represent the inputs of the model, including stop-bar loop detector waveforms and signal timing plans (additinally are present driving behavior parameters and turning-movement counts features), while blue-colored objects represent the outputs of the model, including exit and inflow waveforms at each traffic lane, maximum queue length at each traffic phase, and travel time distribution at each traffic phase. Further details regarding the definition and computation of these MOEs are provided in Section IV. Traffic Data Generation.
The following are the key contributions of our work:
  • MTDT is a multi-tasking digital twin engineered to comprehensively capture the nuanced facets of traffic dynamics as vehicles approach and depart from any signalized intersection. Its outputs encompass fine-grained and lane-wise estimations of two primary tasks: exit and inflow waveform time series, along with accurate and phase-wise estimations of two secondary tasks: maximum queue length waveforms and travel time distributions associated with every phase of movement. The model’s adaptability and responsiveness to the temporospatial intricacies of traffic dynamics are evident in its utilization of input parameters such as relevant driving behavior metrics, signal timing plans, turning movement counts, and signal timing plan parameters, among others.
  • Using our digital twins, it becomes feasible to accurately and reliably predict the impact of loop detector waveforms, signal timing plans, driving behavior parameters, and turning movement counts on lane-wise platoons and related MOEs. Importantly, this predictive capability is achieved at a significantly faster pace compared to microscopic simulators, thanks to the utilization of GPU-based processors. The methodologies proposed in our digital twins incur $O(1)$ sequential computation and can be entirely parallelized using cost-effective GPU computation.
  • The findings of this study hold significance for making data-driven decisions regarding changes in traffic signal timing, formulating traffic strategies, planning transportation systems, and redesigning the infrastructure of roadways and intersections. Our digital twins offer valuable insights and tools for traffic control design and safety management, particularly in the context of smart intersections operating within highly dynamic environments.
We designed several experiments, whose findings aim to showcase the feasibility and efficacy of our multi-tasking approach in designing a deep learning-based and multifaceted digital twin for intersections. We introduce two variants of our digital twin and compare our model with them and with our previously introduced TQAM [15] digital twin. The results of the experiments, underscore the versatility of our proposed digital twin, demonstrating its ability to not only generalize estimations across various intersection typologies and characteristics, but also can successfully simulating ATSPM time series waveforms at lane level of resolution, and accurately estimating several Measures of Effectiveness (MOEs) in both time series and distribution forms at phase (lane-group) of resolution for every approach (movement direction) within an urban intersection.
Figure 1: The inputs and outputs of Deep Multi-Task Learning Digital Twin We simulate the ATSPM time series waveform along with waveform or histogram of several MOEs within 8-phase standard NEMA phasing intersections. MTDT simultaneously estimates downstream exit waveforms of every outflow lane in all directions and upstream inflow waveforms of every inflow lane in all directions with an estimation of instant maximum queue length and travel time distribution associated with each phase of movement for an intersection with arbitrary topology and traffic conditions.
The rest of the paper is organized as follows: Section VI provides an overview of related work, discussing various techniques employed for traffic state estimation. Section III outlines the architecture of our proposed digital twin. Finally, in Section V, we assess and compare the performance of our model through a series of experiments involving various observable and unobservable quantities.
II Problem Definition
This section extends the basic concepts and formal definitions introduced in our previous work to formulate our research problem in this study. For simplicity, the notations and terminologies are considered to be consistent. Detailed descriptions of the notations used in this study are presented in Table 1.
Definition 1. Simulation Record: This term refers to a record of a simulation run represented as a tuple of intersection ID, signal waveform $sig$, turn-movement counts vector $tmc$, driver behavior vector $drv$, lane-wise stop-bar loop detector waveforms $stp$, exit loop detector waveforms $ext$, and inflow loop detector waveforms $inf$, and phase-wise queue length waveforms $ql$, and travel time histograms $tt$.
Definition 2. Simulation Graph: This term refers to a single graph data item constructed using a standardized and generic structure, to represent a single traffic scenario occurring at a certain intersection. In this study, we construct two types of Simulation Graph for each intersection $j$ from each single Simulation Record $s$ each with its own attributes and structure to model exit traffic waveforms ($g_{s,j}=(V,E,X,Z)$) and inflow traffic waveforms ($g^{\prime}_{s,j}=(V^{\prime},E^{\prime},X^{\prime},Z)$). Both of them are directed bipartite graphs that connect each stop-bar loop detector to corresponding exit detectors in the node set $V$ or to inflow detectors in the node set $V^{\prime}$. with edge features $Z=[sig_{j,w},tmc_{j},drv_{j}]$ at each graph defined as the concatenation of the signal timing plan time series, turning-movement counts vectors, and driving behavior parameters.
Notation Description Aggr Size Type $j$ Intersection - 1x9 String $stp$ Waveform at Stop-bar detector 5 sec 48x80 Integer, 0-8 $ext$ Waveform at the exit of the intersection 5 sec 16x80 Integer, 0-8 $inf$ Waveform upstream the intersection 5 sec 12x80 Integer, 0-8 $ql$ Queue length time series 5 sec 8x80 Integer, 0-1200 $tt$ Travel time histogram 5 sec 8x200 Integer, 0-700 $sig$ Signal timing state information 5 sec 8x80 Binary $drv$ Driving behavior parameters 2400 sec 1x9 Float, 0-30 $tmc$ Turning-movement counts ratio 2400 sec 35x35 Float, 0-1 $w$ Size of the prediction window - 1 Integer, 0-150 $s$ Single simulation recorded by SUMO 5 sec - Multiple $A$ Common lane-connectivity for $M_{ext}$ module - 2x22 Binary $A^{\prime}$ Common lane-connectivity for $M_{inf}$ module - 2x72 Binary $v$ multivariate time series for queue length estimation 5 sec 1x7x80 - $v^{\prime}$ multivariate time series for travel time estimation 5 sec 1x7x80 - $M_{ext}$ GNN module for exit waveforms estimation - - torch.nn.Module $M_{inf}$ GNN module for Inflow waveforms estimation - - torch.nn.Module $M_{ql}$ CNN module for queue length estimation - - torch.nn.Module $M_{tt}$ CNN module for travel time estimation - - torch.nn.Module
TABLE 1: Summary of the notations and their definitions.
III Proposed Models
In this section, we present our innovative deep-learning solution for a versatile and multi-purpose traffic intersection digital twin. Our Multi-task Deep Learning Digital Twin (MTDT) is designed to simultaneously grasp various facets of intersection-level traffic flow dynamics. MTDT is jointly optimized to learn four sets of joint tasks. The general architecture of our digital twin is composed of two primary modules that yield base learning representations, and two secondary modules that use the representations learned by primary modules to learn their own tasks and improve the learning process of primary tasks as well by the sharing of their representations. For a comprehensive understanding of MTDT’s architecture, refer to the overview depicted in Figure 2.
Figure 2: Overview of the architecture of MTDT. The architecture of our proposed Multi-componential digital twin framework shows two types of modules are interconnection within the pipeline. GATs modules (highlighted in light orange) primarily execute two tasks, while CNNs modules (depicted in dark blue) are tailored for secondary tasks. During inference mode, red-colored arrows activate to seamlessly integrate the aggregated output of primary modules into the input of secondary modules.
The MTDT pipeline comprises two distinct types of modules: Graph Attention Networks (GATs) [17] modules, including $M_{ext}$ and $M_{inf}$, dedicated to the primary task of estimating loop detector waveforms, and Convolutional Neural Networks (CNNs) [8] modules, including $M_{ql}$ and $M_{tt}$, geared towards secondary tasks such as estimating Measures of Effectiveness (MOEs) like maximum queue length and travel time distribution. More specifically, GATs modules, namely $M_{ext}$ and $M_{inf}$, excel at precise, lane-specific estimations of waveform time series for vehicle platoons approaching and departing the target intersection. They meticulously consider instantaneous temporospatial factors like turning movement counts, driving behavior parameters, and signal timing plans. Subsequently, CNNs modules, such as $M_{ql}$ and $M_{tt}$, utilize the outputs from primary modules alongside instantaneous temporospatial characteristics of the target intersection to estimate MOEs like maximum queue length and travel time distribution. This design ensures adaptability without restrictive assumptions, making the digital twin easily instantiated for various secondary tasks related to traffic dynamics.
The underlying principle of this architecture lies in extracting and leveraging valuable additional information from diverse aspects of the same phenomenon, resulting in enhanced performance and more comprehensive outcomes. Proper normalization techniques tailored to each task are employed to enhance the performance of both primary and secondary tasks collectively.
Detailed information regarding the architecture and learning process within each module is elaborated upon in the subsequent subsections:
III-A Exit Module
Exit subnetwork module $M_{ext}$ is an attentional graph neural network (GAT) with an imputation mechanism, as introduced in [23] to perform the primary task of exit waveform time series $ext$ estimation. Here, the input data is fed in the form of Exit Simulation Graphs with a common connectivity matrix and masked node features to connect a total number of 16 exit loop detectors’ $ext$ waveform time series (for 4 outflow lanes for each approach) to the total number of 48 stop-bar loop detectors’ $stp$ waveform time series as in a general topology of an arbitrary up to 4-way intersection. This module consists of a GATConv [17] message passing layer following a single fully connected linear layer parameterized by weight matrix $B\in\mathbb{R}^{w\times N}$ and activated by a ReLU function to generate and propagate messages of hidden representation inflow lanes to any other connected outflow lanes within the intersection network. The attention mechanism in GATConv is a SINGLE-layer feed-forward neural network, parameterized by a weight vector $a\in\mathbb{R}^{N\times w}$ with ReLU non-linearity.
$a_{ij}=ReLU(a^{T}[Wh_{i}\|Wh_{j}])$
Through this technique, GATConv implicitly specifies and integrates the attention coefficients signifying the importance of each $stp$ waveform time series to each $ext$ waveform.
$\alpha_{ij}=Softamx_{j}(\frac{exp(e_{ij})}{\sum_{k\in N_{i}}exp(e_{ik}))}$
Where $N_{i}$ is the connected (neighboring node) lanes of each lane $i$ in the intersection’s graph representation. A fully connected layer activated by a ReLU function is applied on the output of the GATConv layer to impute missing (unobserved) time series waveforms for exit loop detectors $ext$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N\times w}$.
III-B Inflow Module
Inflow subnetwork module $M_{inf}$ is an attentional graph neural network (GAT) with an imputation mechanism, as a simplified version of multi-layer graph neural network [24] solution introduced in [23]. This modules is responsible to perform the primary task of inflow waveform time series $inf$ estimation. The difference between $M_{ext}$ and $M_{inf}$ is that the same GATs module architecture is adopted to perform a different primary task. Here, GATConv implicitly specifies and integrates the attention coefficients denoting the importance $stp$ waveform time series to each $inf$ waveform using a different graph representation of an intersection that connects each of 12 inflow loop detectors’ $ext$ waveform time series (for 3 inflow lanes for each approach) to the total number of 48 stop-bar loop detectors’ $stp$ waveform time series as in a general topology of an arbitrary up to 4-way intersection. In this setting, this GATs module can impute missing (unobserved) time series waveforms for exit loop detectors $inf$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N^{\prime}\times w}$.
III-C Queue Length Module
Queue Length subnetwork module $M_{ql}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of maximum queue length waveform time series $ql$ estimation. Here, the input data is fed in the form of Multivariate Time series $v_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phase, pair 7 variables including exit waveform time series $ext$ estimated by $M_{ext}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant tempospatial characteristics of urban traffic intersections. In this setting, for simplicity reasons, only the three most effective driving behavior i.e., accel, lcCooperative, minGap in traffic dynamic of intersection as shown in [23]. The architecture of the CNNs module consists of 3 Conv1d layers each followed by a MaxPool1d layer and flattening layer before being fed into a fully connected layer activated by a ReLU non-linearity to simultaneously generate a set of 8 outputs as an estimation to maximum queue length waveform time series for every phase of target standard NEMA eight-phasing intersection.
III-D Travel Time Module
Queue Length subnetwork module $M_{tt}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of travel time histogram $tt$ estimation. Here, the input data is fed in the form of Multivariate Time series $v^{\prime}_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phases, pair 7 variables including inflow waveform time series $inf$ estimated by $M_{inf}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant temporospatial characteristics of urban traffic intersections. In this setting, for simplicity reasons, only the three most effective driving behavior i.e., accel, lcCooperative, minGap in traffic dynamic of intersection as shown in [23]. The architecture of this module remains the same as the $M_{ql}$ module.
III-E Loss Construction
Multi-task Deep Learning Digital Twin (MTDT) is trained on numerous realistic traffic scenarios simulated across eight distinct intersections. It takes stop-bar high-resolution loop detector data, driving behavior parameters, signal timing plan parameters, and turning movement counts as inputs to generate exit and inflow time series waveforms for each lane of movement, alongside maximum queue length time series and travel time distribution for each phase of movement.
During the multi-tasking learning process in MTDT, information is extracted from four sources shown as $M=\{M_{ext},M_{inf},M_{ql},M_{tt}\}$ learning modules. These modules cater to four different yet interconnected tasks denoted as $T=\{ext,inf,ql,tt\}$ each yielding an output $\{y_{t}\}_{t\in T}$ and a computed loss value $\{l_{t}\}_{t\in T}$. The model undergoes simultaneous training across all tasks, with losses applied to facilitate learning across the entire spectrum. The final loss $L=\sum_{t\in T}l_{t}$ is computed as the summation of partial loss values obtained from optimizing all primary and secondary learning tasks.
In the training phase, the learning process for both primary and secondary tasks operates independently while being supervised by log information generated by the SUMO micro-simulator for each target variable. The input information is represented either as Simulation Graphs for GATs modules handling primary tasks or as Multivariate Time Series for CNNs modules addressing secondary tasks. While, in the inference phase, the secondary tasks are facilitated by information derived from primary tasks involved in constructing their inputs (i.e., shown by the activation of red-colored arrows in the Figure 2). The output of secondary modules in inference mode are generated by performing 1D convolutional operation on the resulting 1D Multivariate Time Series. It’s worth noting that this study employs a simplified design to illustrate the benefit of this generic approach, more complex architectures with improved feature learning might lead to better results.
To optimize the generalization of the optimization scheme within our multi-task learning paradigm, we employ appropriate normalization techniques tailored to each task to ensure equal weighting of partial loss values computed by different tasks. $M_{ext}$, $M_{inf}$, and $M_{ql}$ is optimized using the mean squared error (squared L2 norm) loss metric while $M_{tt}$ is optimized by cross-entropy loss function with the number classes equal to the number of aggregated bins in the original histogram of the distribution. The normalization technique used for $y_{ext}$ and $y_{inf}$ and $y_{tt}$ is min-max normalization, while $y_{ql}$ is log scaled. For each normalization and scaling formula, a counterpart denormalization function is utilized to transform $\{\hat{y}_{t}\}_{t\in T}$ back to the original scale before computing error values for estimations.
$\hat{y}_{t}=Denorm_{t}(M_{t}(Norm_{t}(X_{t})))$
The model is trained on the train portion, and a grid search algorithm is employed to fine-tune regularization hyperparameters. The model is evaluated on to the test partition and the error values are reported for the best-performing model across 100 epochs of training.
In this study, the secondary sub-network modules are designed with as few parameters as possible and low computational overhead. This design ensures easy integration into the primary block, enhancing the efficiency of multi-task learning within the resulting model. While we deliberately maintain a simplistic architecture to showcase the advantages of this learning paradigm in our application, introducing subtle additional complexities could potentially yield improved results through enhanced feature learning capabilities.
IV Traffic Data Generation
Our dataset is based on over 400,000 simulation-hours of SUMO data. A 9-intersection urban corridor located in a large metropolitan region in the United States of America was chosen. The dataset has been described in [23] in detail. We reiterate the important points here.
We use real-world ATSPM (loop detector) data with sparse WEJO (GPS) probe trajectory data sources to generate an approximate Origin-Destination probability matrix. This provides insight into the probability of a vehicle originating from a specific location reaching various valid destinations. SUMO tool od2trips is used to generate route files for the simulation.
Building upon the configuration parameters and the range of variation detailed in [23], we augment our data generation process with additional processing. This enhancement enables the collection of several Measures of Effectiveness (MOEs) corresponding to the state of traffic flow across various simulated counterfactual traffic scenarios.
For signal timing plans, we use Ring-and-Barrier operation with a common cycle length. Parameters such as the minimum/maximum green times, yellow times, and red times for the intersections are based on signal timing sheets. However, we do vary the common cycle lengths and barrier times. A random common cycle length is chosen between 120 seconds and 240 seconds. Once this is selected, random signal cycle offsets are chosen for each intersection separately. A random barrier time is chosen. Driving behavior parameters (such as acceleration, deceleration, emergency braking, minimum gap between vehicles, headway, lane-changing parameters, etc.) are also varied. Finally with some additional processing, we get the outputs of several Measures of Effectiveness(MOEs) to measure the state of traffic flow generated at counterfactual scenarios.
Queue Length. The maximum queue length time series for each approach or traffic phase represents the instant magnitude of the maximum queue length in the unit of meter aggregated at 5-second bucket resolution. In order to get the variation of queue length over time, we take the maximum queue length seen for a group of lanes associated with each phase of movement at a 5-second resolution. This is obtained directly by parsing SUMO log files and tracking queue lengths on a per-second basis. For the through-lanes (i.e. lanes serving the even number phases 2/4/6/8), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the through-lanes and (b) 2-hop lanes w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. Broadly, this number gives the maximum possible queue length that any of the lanes belonging to the through-lanes, might experience. For the left-lanes (i.e. lanes serving the odd number phases 1/3/5/7), we take the maximum queue lengths seen in 5 seconds across all (a) the 1-hop lanes serving the left-lanes and (b) the left-most 2-hop lane w.r.t. the intersection. We then sum these two maximum queue lengths (a) and (b) to get the overall queue length. The reason for choosing only the left-most 2-hop lane for (b) is that it is assumed that many vehicles that wish to take a left-turn, would have begun changing into the left-most lane in anticipation of the start of the left-turn buffer. Thus, the maximum queue length should take into account the left-most lane.
Intersection Travel Time. The travel time histogram for each approach or traffic phase represents the frequency of vehicles that completed that traffic movement journey at the intersection groping into intervals of 5 seconds within a travel time range of 0 to 1000 seconds. This is obtained from SUMO logs, by processing its trajectory data at 1 second resolution. A vehicle is said to be within the proximity of the intersection if the vehicle is on a lane that is 1-hop or 2-hop away. The travel time here tracked is the amount of time the vehicle spends in the proximity of the intersection till it exits the intersection. For each vehicle, its travel time with respect to the intersections it crosses, is extracted. These travel times are collected for each intersection and each group of lanes associated with a phase of movement and a histogram is created. The histogram varies from 0 to 1000 seconds, in buckets of 5 seconds each. Each bucket counts how many vehicles had that travel time. Eg. a count of 7 vehicles in the bucket 45-50 s along phase 2 of intersection 5, means that 7 vehicles took between 45 and 50 seconds to cross the proximity of that intersection.
V Experimental Results
In this section, we assess the performance of our proposed multi-tasking digital twin, referred to as MTDT. This MTDT model undergoes training on a dataset comprising 22,000 Simulation Records meticulously gathered from a diverse range of eight intersections. Additionally, we introduce two variants for comparison: MTDT-Single, trained on a dataset of 19,000 Simulation Records from a single intersection, and MTDT-MOE, a modified version of MTDT that excludes primary modules $M_{ext}$ and $M_{inf}$ from the pipeline, resulting in the output of only two sets of MOEs variables, namely $ql$ and $tt$, instead of the full set of variables. Lastly, we include TQAM [15] a RNN Auto-encoder as a baseline deep learning-based digital twin previously introduced in our research for the estimation of maximum queue length.
We employ widely used metrics to assess our model’s performance, including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Normalized Root Mean Squared Error (NRMSE). For travel time distribution estimation, we dissect error measures for the 60th, 75th, 85th, and 90th percentiles separately. This breakdown allows us to scrutinize the model’s performance across various traffic flow dynamics within each percentile. When estimating queue length waveform time series, we opt for NRMSE over RMSE to ensure error values are relative to the scale of data. Additionally, to mitigate scale-related biases, we segment our dataset based on the magnitude of the maximum queue length logged by each Simulation Records and report error values for each group individually.
Table. 2 shows our proposed digital has the potential to effectively learn and generalize intersection-level traffic simulation to any arbitrary intersection when it is exposed for an adequate number of epochs to the different intersections. MTDT-Single and MTDT show similar results for inflow and exit waveform time series and travel time distribution estimation. Although MTDT-Single outperforms in the estimation of the maximum queue length, the comparison, in general, implies that our digital twin is able to effectively generalize the joint estimation of traffic flow and MOE estimation at the intersection level regardless of its lane design and configuration (topology). The benefit of multi-task learning can be proved by comparing the performance of MTDT-MOE and MTDT-Single. The performance of secondary tasks is supported by the outputs of the primary task, resulting in better performance in MTDT-Single compared to MTDT-MOE.
The minimal marginal increase in error values for all variable remains minimal relative to the aggregation level indicates consistent reliability and robustness in the model’s performance across varying data resolutions. We also note the model exhibits enhanced performance in estimating travel time for lower quantiles of the data distribution while showing a slight improvement, particularly at the 90th quantile. This suggests that our designed digital twin is also particularly adept at capturing relatively rare or extreme patterns.
Aggregation MTDT-Single MTDT MTDT-MOE Maximum Queue Length Estimation $MAE|RMSE$ $MAE|RMSE$ $MAE|RMSE$ 5-second buckets $4.8084|0.0885$ 9.6114||||0.04621 $9.0321|0.0297$ 10-second buckets $9.3787|0.08410$ $18.9512|0.08118$ $17.8554|0.0513$ 15-second buckets $13.8788|0.0973$ $27.9748|0.1012$ $26.4127|0.0674$ 20-second buckets $18.2466|0.0907$ $37.1869|0.1003$ $35.1794|0.0688$ Travel Time Distribution Estimation $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ 60th percentile $5.50946|17.744|\mp 34.7782$ $5.5332|17.7440|\mp 34.7782$ $4.8024|15.3052|\mp 29.9981$ 75th percentile $5.7687|17.7987|\mp 34.8854$ $5.7592|18.3542|\mp 35.9742$ $4.8387|15.2256|\mp 29.8421$ 85th percentile $4.7090|13.0162|\mp 25.5117$ $5.7298|18.1125|\mp 35.5005$ $4.6844|14.8253|\mp 29.0575$ 90th percentile $4.1331|11.7298|\mp 22.9904$ $5.0685|16.3674|\mp 32.0801$ $3.7703|12.3752|\mp 24.2553$ Exit Waveform Estimation $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ 5-second buckets $0.0979|0.2995|\mp 0.5870$ $0.1020|0.3189|\mp 0.6250$ 10-second buckets $0.1758|0.5083|\mp 0.9962$ $0.1844|0.5575|\mp 1.0927$ 15-second buckets $0.2477|0.6977|\mp 1.3674$ $0.2607|0.7761|\mp 1.5211$ 20-second buckets $0.3126|0.8683|\mp 1.7018$ $0.3326|0.9775|\mp 0.9775$ Inflow Waveform Estimation $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ $MAE|RMSE|95\%CI$ 5-second buckets $0.1145|0.3456|\mp 0.6773$ $0.1284|0.3785|\mp.7418$ 10-second buckets $0.2119|0.6005|\mp 1.1769$ $0.2385|0.6729|\mp 1.3188$ 15-second buckets $0.3018|0.8287|\mp 1.6242$ $0.3397|0.9387|\mp 1.8398$ 20-second buckets $0.3837|1.0340|\mp 2.0266$ $0.4367|1.1844|\mp 2.3214$
TABLE 2: Performance of MTDT. The performance of our proposed digital twin is assessed across various aggregation resolutions of maximum queue length and at different quantiles of the travel time distribution. MTDT is compared to its variant (MTDT-Single) when trained on a single intersection, and MTDT-Single is further compared to MTDT-MOE, where GAT modules of primary tasks are ablated from the pipeline. This comparison is conducted in terms of NRMSE, RMSE and MAE, measured in vehicle-per-bucket for Exit and Inflow waveforms, meter-per-bucket for maximum queue length, and vehicle-per-quantile for travel time distribution. Moreover, the level of confidence for estimating the relevant variable across various resolution levels are shown as 95% confidence intervals (CI).
For the multifaceted and fine-grained results, our designed multi-tasking simulation stands as unique to its counterparts, thus not directly comparable to existing baselines. However, we attempt to contrast our findings with those obtained by TQAM designed to make an accurate estimation of queue length estimation in a specific approach of a specific topology of target intersection. To achieve this, we partition our data into the same queue length buckets as reported for this model and compare error values in terms of (percentual) Mean Absolute Error (MAE) in Table 3. The partitioning encompasses two groups for each of the low (L), medium (M), and high (H) ranges of maximum queue length. These groups, ranging from L1 to H2, encompass 28%, 34%, 19%, 11%, 5%, and 3% of data samples, respectively. Our observations indicate that both MTDT and MTDT-Single exhibit comparable performance to TQAM. Specifically, MTDT outperforms TQAM in certain partitions of data where very short (0-40 meters) or long (above 200 meters) queue length tails are recorded, while MTDT-Single consistently performs equivalently or better than TQAM across nearly all partitions.
Max. Queue
Length (meter):
L1
0-40 m
L2
40-80 m
M1
80-120 m
M2
120-160 m
H1
160-200 m
H2
200+ m
MTDT
3.02 (7.5%) 6.00 (7.5%) 8.65 (7.2%) 11.34 (7.0%) 16.46 (8.2 %) 15.26 (1.5%)
MTDT-Single
2.94 (7.3 %) 5.42 (6.7 %) 7.27 (6.0 %) 9.16 (5.7 %) 11.05 (5.5 %) 9.80 (0.9 %)
TQAM 2.90 (7.3%) 5.50 (6.9%) 7.00 (5.8%) 9.00 (5.6%) 11.60 (5.8%) 18.10 (1.8%)
TABLE 3: Evaluation at various maximum queue length ranges A comparative analysis is conducted between MTDT, MTDT-Single, and TQAM for queue length waveform time series estimation across different queue length buckets. The assessment is based on MAE and the percentage of MAE relative to the upper range within each bucket, measured in meters per 5-second time resolution.
We further devise an experiment to assess the performance of our proposed digital twin under varying percentages of green time allocation to the major direction of the intersection. In this experimental setup, we focus on Phase 2/6 in the corridor-through phase, serving the major direction of traffic movement for all intersections. We define low, medium, and high percentages of this green time as 45-50%, 60-75%, and 75-90% of the total duration of the simulation scenario. However, the individual cycle lengths may vary due to actuated ring-and-barrier operation responding to the cycle-wise incoming traffic. After categorizing simulation records into these three groups, we evaluated the performance of MTDT and MTDT-Single across all output variables, as presented in Table 4. The findings indicate that the performance of our digital twin is fairly consistent. It excels in queue length estimation under scenarios with shorter green time allocated to the major direction, whereas travel time estimation performs better with a larger green time allocation.
Green Time (percentage):
Low
45-60 %
Medium
60-75 %
High
75-90 %
Task Error Measure
MTDT-Single
Exit Waveform
MAE
RMSE
0.0106
0.3556
0.0963
0.3121
0.0936
0.3439
Inflow Waveform
MAE
RMSE
0.1518
0.1199
0.3632
0.320
0.4113
0.3430
Maximum Queue Length
MAE
NRMSE
6.0780
0.2071
5.9700
0.1392
4.5635
0.0778
Travel Time Distribution
MAE
RMSE
1.4173
4.9942
2.8246
9.7389
3.1870
10.827
MTDT
Exit Waveform
MAE
RMSE
0.0905
0.3288
0.09108
0.0800
0.0926
0.3052
Inflow Waveform
MAE
RMSE
0.1204
0.4052
0.1151
0.3811
0.1101
0.3438
Maximum Queue Length
MAE
NRMSE
12.105
0.4243
11.946
0.2780
8.6872
0.1557
Travel Time Distribution
MAE
RMSE
0.0915
8.4421
2.3671
9.0937
3.0768
10.735
TABLE 4: Evaluation at various green time allocations. The performance of the model is evaluated at different plit of data based on corridor-through green time percentage buckets. The error values are reported based on NRMSE, MSE and MAE values measured in vehicle per 5-second bucket for Exit and Inflow waveforms, meter per 5-second bucket for maximum queue length, and vehicle per 60th percentile for travel time distribution.
VI Related Work
Existing methods of traffic flow modeling in the past literature can be categorized into linear, non-linear, machine learning, and hybrid models. With advancements in high-resolution induction loop detectors in intelligent transportation systems (ITSs), a large number of deep neural network models have been proposed for real-time and short-term traffic flow prediction raised by industries, academia, and government. In the realm of linear models, [1] proposed a localized (dynamic weight matrix) space-time autoregressive (LSTAR) model for real-time traffic prediction, [4] developed a Kalman filter (KF) for predicting real-time traffic flow of connected vehicles at urban arterials, [12] combines KF and adaptive weight allocation, and [14] formulates a consensus filter-based distributed variational Bayesian (CFBDVB) algorithm for density approximation of traffic flow and average traffic speed in a freeway.
In the realm of non-linear models, multi-kernel SVM [10] and weighted k-nearest neighbor [21] with MapReduce algorithm on Hadoop are introduced. When it comes to hybrid models, ensemble empirical mode decomposition with adaptive noise [16] and bidirectional RNN combined with road crossing vector coding [26] can be mentioned. Machine learning models in this line of research are deep learning frameworks based on the temporal convolutional network (TCN) [27], parallel computing of deep belief networks (DBNs) [25], deep characteristic learning using DBN with the multi-objective particle swarm optimization (PSO) algorithm [9]. [3] and [11] developed a deep hybrid neural network combined with CNN and LSTM using trajectory data for better accuracy and robustness of forecasting performance, while [paper[1]] combined GCN and attention mechanism for lane-wise and topology invariant traffic flow estimation.
Some literature integrates operation performance forecasting in traffic signal control of intersections and networks. For example, [13] makes use of graph mining and trajectory data mining methods to provide interactive evaluation and exploration of the various MOEs, while [2] proposed a safety-enhanced residual reinforcement learning method (SafeLight) to integrate safety conditions in traffic signal control by integrating knowledge using multi-objective loss function and reward shaping, and [6] proposed a metaheuristic-based method for intelligent traffic control using Genetic Algorithm (GA) and differential evolution (DE) to improve the level of service (LOS) i.e. delay optimization of intersection through optimizing the signal timing plan.
VII Conclusions and Future Work
In this paper, we demonstrate the effectiveness of multi-tasking as a learning paradigm for deep learning-based digital twins. We introduce a Multi-Task Deep Learning Digital Twin, so-called MTDT, capable of successfully performing diverse albeit related traffic simulation tasks related to an urban intersection with an arbitrary topology and characteristics. Our findings highlight the benefits of the multi-task learning approach in the design on our digital twin, not only for secondary tasks whose inputs rely on the output of primary tasks but also for primary tasks to generalize and enhance their performance through shared learned representations and joint optimization. Furthermore, we design two experiments to specifically illustrate how the model’s performance varies across different partitions of data when simulation scenarios are segmented based on signal timing plans or maximum recorded queue length. The results underscore the consistency and robustness of our approach across different experiments. Moreover, we discuss the lack of baseline due to the uniqueness and comprehensiveness of our approach. Although, we show it still outperforms in individual tasks compared to some previously introduced digital twins specialized for that specific tasks.
Looking ahead, there are several aspects of future work in this research that we briefly discuss in the following.
In this approach, we use lane-wide inflow time series waveforms $inf$ estimated by the graph-based digital twin $G_{inf}$ as vehicles approach the target intersection from various directions, we construct the Origin-Destination (O-D) matrix. Subsequently, we employ the od2trips tool of SUMO to generate route files and evaluate each traffic scenario using a variety of Measures of Effectiveness (MOEs). An Optimum signal timing plan can be viewed as the most favorable trade-off among desired MOEs, considering a range of traffic conditions such as signal timing plans and driving behaviors. We consider this topic a significant avenue for our future research.
MTDT can serve leveraged as a valuable tool to assist traffic decision-makers in comprehending the potential impact of alterations to signal timing plans and proposed changes to road geometry. Additionally, it can be seamlessly integrated into signal timing optimization frameworks, including those based on reinforcement learning and other methodologies.
Moreover, this approach can be expanded from isolated intersections to corridors and networks, paving the way for traffic signal timing plan optimization, which will be a focal point of our future endeavors. Moreover, our model holds promise for integration into MapReduce frameworks, facilitating distributed training and parallel prediction of traffic flow.
VIII Acknowledgments
The work was supported in part by NSF CNS 1922782. The opinions, findings, and conclusions expressed in this publication are those of the authors and not necessarily those of NSF. The authors also acknowledge the University of Florida Research Computing for providing computational resources and support that have contributed to the research results reported in this publication.